OpenStudy (anonymous):

A map has a scale of 3 km to 2 inches. Which ratio of measures forms a proportion with the map's scale of km to inches? A- 7/8 B- 8/12 C- 6/5 D- 15/10 Please help need answer right away!

OpenStudy (mathstudent55):

The map's scale is the ratio 3/2 Now look at each choice and reduce to lowest terms and see which one equals 3/2 A - 7/8 is in lowest terms, not equal to 3/2 B - 8/12 = 2/3, not equal to 3/2 C - 6/5 is in lowest terms, not equal to 3/2 D - 15/10 = 3/2 which is what the map has

OpenStudy (anonymous):

Oh Okay, Thank Youu!

OpenStudy (mathstudent55):

You're welcome.